# Visual Agents as Fast and Slow Thinkers

### v2

@@ 4 Experiment
We utilize eight popular benchmarks to evaluate our framework FaST comprehensively, categorized into general visual question answering (VQA) datasets and multimodal benchmarks. The VQA benchmarks include VQA-v2 [106], GQA [107], ScienceQA [108], and TextVQA [27] which focus on optical character recognition. For multimodal benchmarks evaluation, we use the hallucination benchmark POPE [109], along with comprehensive benchmarks such as MME [28], MM-Vet [110], and SEED [111]. We compare our model with the baseline LLaVA-v1.5 [112], and other multimodal large language models. To thoroughly assess our model’s understanding of pixel-level instances, we evaluate its performance on referring segmentation and grounding benchmarks, including refCOCO [105], refCOCO+ [105], and refCOCOg [99]. Further, to examine the model’s reasoning capabilities on FaST framework, we consider the Reasoning Segmentation benchmark [29].
@@ 4.1.1 Experiments on VQA and Multimodal Benchmarks
Training. In developing the Switch Adapter, we employed the LLaVA-v1.5 [26] framework, conforming strictly to its established training protocols. We incorporated negative samples from $V^{*}$ [19] with contextual cues to enhance system switching capability to amplify multimodal inferential and world knowledge. This augmented dataset was combined with LLaVA-v1.5’s supervised dataset and trained for one epoch. For the Proposal Adapter, we augmented the LLaVA-v1.5 dataset with region-specific bounding boxes based on contextual cues and queries, then fine-tuned for one epoch to optimize proposal generation. The Segmentation Adapter utilized the LISA [29] architecture integrated with the LLaVA-v1.5, employing SAM as the mask decoder. The adapter was fine-tuned using the same datasets as Lisa, including semantic segmentation, referring segmentation, and reasoning segmentation. This fine-tuning process involved 10,000 steps to improve the model’s segmentation capabilities. Throughout developing the Switch Adapter, Proposal Adapter, and Segmentation Adapter, we employed the LoRA (Low-Rank Adaptation) technique [113]. By leveraging LoRA, we introduce minimal additional parameters while preserving the original multimodal large language model’s architecture and efficiency. All experiments used 8 NVIDIA TESLA A100-80GB GPUs.
@@
Results. As depicted in Table 1, FaST demonstrates superior performance across multiple VQA datasets and multimodal benchmarks when compared to established methods. Notably, FaST surpasses the LLaVA-v1.5 model, with marked improvements in performance across all datasets. Specifically, in the VQA datasets, our model achieves an enhanced accuracy of 2.3% in $VQA^{v2}$, 1.8% in GQA and 2.5% in $VQA^{T}$ comparing to LLaVA-1.5. Similarly, in the multimodal benchmarks, FaST enhances performance notably with an increment of 6.7 in the MME, 1.5 in the SEED and 0.5 in the MM-Vet score, underlining its effectiveness in handling diverse domains. This consistent performance underscores the robustness of FaST, especially in complex visual and textual tasks.
  VQA Datasets Multimodal Benchmarks Method LLM $VQA^{v2}$ GQA $VQA^{T}$ $SQA^{I}$ POPE MME SEED MM-Vet BLIP-2[ICML23] [36] Vicuna-13B 65.0 32.3 42.5 61.0 85.3 1293.8 46.4 22.4 InstructBLIP[NeurIPS24] [37] Vicuna-13B - 49.5 50.7 63.1 78.9 1212.8 53.4 25.6 Qwen-VL-Chat[arXiv23] [41] Qwen-7B 78.2 57.5 61.5 68.2 - 1487.5 58.2 - mPLUG-Owl2[CVPR24] [45] LLaMA-7B 79.4 56.1 58.2 68.7 - 1450.2 61.6 36.2 Monkey[CVPR24] [114] Qwen-7B 80.3 60.7 - 69.4 67.6 - - - LLaVA-v1.5[CVPR24] [26] Vicuna-7B 78.5 62.0 58.2 66.8 85.9 1510.7 58.6 30.5 FaST (Ours) Vicuna-7B 80.8 63.8 60.7 68.9 86.4 1517.4 60.1 31.0 $\Delta$ (vs LLaVA-v1.5) Vicuna-7B +2.3 +1.8 + 2.5 +2.1 +0.4 +6.7 + 1.5 + 0.5
@@ 4.1.2 Experiments on Referring and Reasoning Segmentation
Metric. Following prior research on segmentation [105, 104], two evaluation metrics are employed: Generalized Intersection over Union ($GIoU$) and complete Intersection over Union ($CIoU$).
@@
$GIoU$. The $GIoU$ is computed as the average per image $IoU$, where the $IoU$ is calculated for each individual image, and then the average is taken across all images in the dataset. This metric provides a balanced assessment by treating all images equally, regardless of their size or the number of objects.
@@
 LAVT[CVPR22] [115] 72.7 62.1 61.2 - -
OVSeg[CVPR23] [116] - - - 28.5 18.6
GRES[CVPR23] [117] 73.8 66.0 65.0 22.4 19.9
X-Decoder[CVPR23] [118] - - 64.6 22.6 17.9
SEEM[NeurIPS24] [119] - - 65.7 25.5 21.2
LISA-7B[CVPR24]  [29] 74.1 62.4 66.4 44.4 46.0
@@
Results. Table 2 presents the results of referring and reasoning segmentation benchmarks. FaST compares favorably to recent visual agents like LISA-7B. Notably, FaST outperforms LISA-7B on the refCOCO+ and refCOCOg benchmarks by a substantial margin of 2.0% and 0.6% $CIoU$, respectively. For the more challenging reasoning segmentation task, FaST achieves a significant improvement over LISA-7B, with 3.2% $CIoU$ and 2.7% $GIoU$ gain, demonstrating its efficacy across diverse segmentation benchmarks and showcasing its robustness in handling complex visual reasoning queries.
@@ 4.2 Analysis of System Switching Adapter
Our study investigates the efficacy of the switch adapter mechanism in balancing accuracy and computational efficiency. As depicted in Fig.4, our analysis illustrates the system’s adeptness in discerning between the System 1 and System 2 cognitive modes triggered by query complexity. Notably, the switch adapter activates the System 1 mode in 41.8% and 64.4% of cases for the MME and GQA datasets, respectively. Notably, responses generated under the System 1 mode exhibit significantly higher accuracy rates (78.8% for MME and 64.7% for GQA) compared to those under System 2 mode (52.2% for MME and 56.8% for GQA), highlighting the adapter’s ability to differentiate query complexities and optimize task performance accordingly, emphasizing the importance of maintaining a fast thinking mode for prompt and confident responses.
@@
Table 3 offers valuable insights into the runtime comparison of different system configurations. System 1 Only operates with just a switch adapter, with the slow thinking process disabled. For System 2 Only, we force to construct chain of evidence for every single query. Notably, System 1 only operates efficiently, requiring only one-time inference. In contrast, System 2 only demands significantly more computational resources. Meanwhile, our FaST offers a strategic advantage by optimizing the balance between computational efficiency and performance. It operates approximately 31% faster than System 2 Only for MME and about 50% faster for GQA, while maintaining comparable performance outcomes. FaST enhances the overall efficiency of cognitive processing tasks, offering a pragmatic approach to optimizing computational resources.
@@ 4.3 Ablation Study
Key Component Analysis. We investigate the core elements of our novel framework FaST, focusing on the proposal adapter for contextual region localization and the seg adapter for pixel-level mask segmentation. To establish a Baseline, we deploy a model devoid of proposal adapter and seg adapter, relying solely on a switch adapter to supply missing objects and context clues. As shown in Table 5, after integrating the proposal adapter, seg adapter, or both, we observe progressive and significant enhancements in performance across various evaluation metrics (62.1% $\rightarrow$ 63.8% accuracy on $VQA^{v2}$). These findings underscore the pivotal roles played by the proposal and seg adapters in augmenting the model’s capabilities, thereby affirming their significance within the FaST framework.
Furthermore, we analyze the output components of our switch adapter in FaST, focusing on including missing objects and context clues. We construct Baseline* model which operates without these enhancements. As illustrated in Table 5, adding the missing objects component shows a noticeable improvement in both MME and refCOCOg scores. Similarly, the integration of context clues also leads to further performance gains, albeit not as significant as missing objects. When both components are integrated, the model demonstrates the best result metrics. These results highlight the significant impact of incorporating missing objects and context clues on the model’s overall effectiveness.
@@ 4.4 Qualitative Comparisons of FaST
In Fig.5, we present qualitative comparisons to illustrate the enhancements introduced by FaST. The top row above the dotted line shows the results for the VQA task, comparing FaST with LLaVA-v1.5. LLaVA-v1.5 fails to identify the correct focusing areas, resulting in incorrect responses. In contrast, FaST leverages the extensive world knowledge of LLMs to construct a chain of evidence ($e.g.$, it can identify a woman on the street or a monitor on the roof). Then FaST determines the appropriate focus area with object-level pixel masks for an accurate and deliberate answer. The bottom row under the dotted line displays segmentation results compared with LISA-7B. LISA-7B exhibits confusion in accurately distinguishing items that are comparatively smaller and require more extensive reasoning to discern. Conversely, FaST effectively isolates the relevant objects with greater granularity and accuracy. This demonstrates the superior capability of FaST in both VQA and segmentation tasks.

### v3

@@ 3 Experiment
We utilize eight popular benchmarks to evaluate our framework FaST comprehensively, categorized into general visual question answering (VQA) datasets and multimodal benchmarks. The VQA benchmarks include VQA-v2 (Goyal et al., 2017), GQA (Hudson & Manning, 2019), ScienceQA (Lu et al., 2022), and TextVQA (Singh et al., 2019) which focus on optical character recognition. For multimodal benchmarks evaluation, we use the hallucination benchmark POPE (Li et al., 2023c), along with comprehensive benchmarks such as MME (Fu et al., 2024), MM-Vet (Yu et al., 2024), and SEED (Li et al., 2024). We compare our model with the baseline LLaVA-v1.5 (Liu et al., 2023a), and other multimodal large language models. To thoroughly assess our model’s understanding of pixel-level instances, we evaluate its performance on referring segmentation and grounding benchmarks, including refCOCO (Kazemzadeh et al., 2014), refCOCO+ (Kazemzadeh et al., 2014), and refCOCOg (Caesar et al., 2018). Further, to examine the model’s reasoning capabilities on FaST framework, we consider the Reasoning Segmentation benchmark (Lai et al., 2024).
@@ 3.1.1 Experiments on VQA and Multimodal Benchmarks
Training. In developing the Switch Adapter, we employed the LLaVA-v1.5 (Liu et al., 2024a) framework, conforming strictly to its established training protocols. We incorporated negative samples from $V^{*}$ (Wu & Xie, 2024) with contextual cues to enhance system switching capability to amplify multimodal inferential and world knowledge. This augmented dataset was combined with LLaVA-v1.5’s supervised dataset and trained for one epoch. For the Proposal Adapter, we augmented the LLaVA-v1.5 dataset with region-specific bounding boxes based on contextual cues and queries, then fine-tuned for one epoch to optimize proposal generation. The Segmentation Adapter utilized the LISA (Lai et al., 2024) architecture integrated with the LLaVA-v1.5, employing SAM as the mask decoder. The adapter was fine-tuned using the same datasets as Lisa, including semantic segmentation, referring segmentation, and reasoning segmentation. This fine-tuning process involved 10,000 steps to improve the model’s segmentation capabilities. Throughout developing the Switch Adapter, Proposal Adapter, and Segmentation Adapter, we employed the LoRA (Low-Rank Adaptation) technique (Hu et al., 2022). By leveraging LoRA, we introduce minimal additional parameters while preserving the original multimodal large language model’s architecture and efficiency. All experiments used 8 NVIDIA TESLA A100-80GB GPUs.
@@
Figure 4: The Comparison with CoVLM and CogCoM. These models use the more powerful vison encoder.
@@
Results. As depicted in Table 1, FaST demonstrates superior performance across multiple VQA datasets and multimodal benchmarks when compared to established methods. To ensure fairness in comparison, all methods in Table 1 share the same visual encoder: basic CLIP (Radford et al., 2021). Remarkably, FaST consistently surpasses the LLaVA-v1.5 model, achieving significant improvements in performance across all evaluated datasets. Specifically, in VQA datasets, our model outperforms LLaVA-v1.5 by 2.3% in $VQA^{v2}$, 1.8% in GQA, and 2.5% in $VQA^{T}$. Additionally, FaST excels in multimodal benchmarks, with notable increases of 6.7 in the MME score, 1.5 in the SEED score, and 0.5 in the MM-Vet score, highlighting its versatility and effectiveness in handling a broad range of domains. These results underscore the robustness of FaST, particularly in tackling complex visual and textual tasks. Moreover, Fig. 4 showcases a direct comparison between FaST, CoVLM (Wang et al., 2023b), and CogCoM (Qi et al., 2024), both of which employ the more powerful EVA2-CLIP-E (Sun et al., 2023) model as their visual encoder. As expected, these models exhibit stronger performance due to their enhanced encoder. To align with this, we replaced our original vision encoder with EVA2-CLIP-E, which resulted in further improved performance, ensuring a more rigorous and fair comparison with state-of-the-art methods. This two-tiered comparison—first with basic CLIP and then with the more advanced EVA2-CLIP-E—provides a balanced and comprehensive evaluation of FaST against leading approaches, reinforcing its effectiveness in diverse and challenging tasks.
  VQA Datasets Multimodal Benchmarks Method LLM $VQA^{v2}$ GQA $VQA^{T}$ $SQA^{I}$ POPE MME SEED MM-Vet BLIP-2[ICML23] Vicuna-13B 65.0 32.3 42.5 61.0 85.3 1293.8 46.4 22.4 InstructBLIP[NeurIPS24] Vicuna-13B - 49.5 50.7 63.1 78.9 1212.8 53.4 25.6 Qwen-VL-Chat[arXiv23] Qwen-7B 78.2 57.5 61.5 68.2 - 1487.5 58.2 - mPLUG-Owl2[CVPR24] LLaMA-7B 79.4 56.1 58.2 68.7 - 1450.2 61.6 36.2 Monkey[CVPR24] Qwen-7B 80.3 60.7 - 69.4 67.6 - - - LLaVA-v1.5[CVPR24] Vicuna-7B 78.5 62.0 58.2 66.8 85.9 1510.7 58.6 30.5 Chain of Spot [arXiv24] Vicuna-7B 80.7 63.7 60.9 68.2 86.4 1501.1 59.7 30.8 V*[CVPR24] Vicuna-7B - - - - 82.4 1128.9 41.7 27.7 Visual CoT[arXiv24] Vicuna-7B - 63.1 77.5 - - - - - FaST (Ours) Vicuna-7B 80.8 63.8 60.7 68.9 86.4 1517.4 60.1 31.0 $\Delta$ (vs LLaVA-v1.5) Vicuna-7B +2.3 +1.8 + 2.5 +2.1 +0.4 +6.7 + 1.5 + 0.5
@@ 3.1.2 Experiments on Referring and Reasoning Segmentation
Metric. Following prior research on segmentation (Kazemzadeh et al., 2014; Mao et al., 2016), two evaluation metrics are employed: Generalized Intersection over Union ($GIoU$) and complete Intersection over Union ($CIoU$).
@@
$GIoU$. The $GIoU$ is computed as the average per image $IoU$, where the $IoU$ is calculated for each image, and then the average is taken across all images in the dataset. This metric provides a balanced assessment by treating all images equally, regardless of their size or the number of objects.
@@
LAVT[CVPR22] 72.7 62.1 61.2 - -
OVSeg[CVPR23] - - - 28.5 18.6
GRES[CVPR23] 73.8 66.0 65.0 22.4 19.9
X-Decoder[CVPR23] - - 64.6 22.6 17.9
SEEM[NeurIPS24] - - 65.7 25.5 21.2
LISA-7B[CVPR24] 74.1 62.4 66.4 44.4 46.0
@@
Results. Table 2 illustrates the performance of FaST compared to recent visual agents like LISA on referring and reasoning segmentation benchmarks. FaST notably outperforms LISA-7B on the refCOCO+ and refCOCOg benchmarks by 2.0% and 0.6% $CIoU$, respectively. For the more complex reasoning segmentation task, FaST shows even stronger results, with a 3.2% $CIoU$ gain and a 2.7% $GIoU$ improvement over LISA. The results highlight FaST’s superior performance and its robustness in handling both straightforward and complex visual reasoning segmentation benchmarks.
@@ 3.2 Analysis of System Switching Adapter
Our study investigates the efficacy of the switch adapter mechanism in balancing accuracy and computational efficiency. As depicted in Fig.5, our analysis illustrates the system’s adeptness in discerning between the System 1 and System 2 cognitive modes triggered by query complexity. For queries requiring System 2 reasoning, the adapter dynamically combines System 1 reasoning for simpler subcomponents with System 2 reasoning for more complex aspects. Consequently, the reported accuracy rates under System 2 mode (52.2% for MME and 56.8% for GQA) reflect a combination of reasoning outcomes, emphasizing the adapter’s ability to differentiate query complexities and optimize task performance accordingly. This highlights the importance of maintaining System 1 reasoning for prompt and confident responses while effectively utilizing System 2 reasoning for complex problem-solving.
@@
Table 3 compares runtime across system configurations. System 1 Only, using a switch adapter, operates efficiently with one-time inference, while System 2 Only, which constructs a chain of evidence for every query, is significantly more resource-intensive. In contrast, FaST balances efficiency and performance, running 31% faster than System 2 Only on MME and 50% faster on GQA, with comparable results. This highlights FaST’s ability to optimize cognitive task processing while conserving computational resources.
@@ 3.3 Ablation Study
Key Component Analysis. We undertake a detailed investigation into the core elements of our novel framework, FaST, with particular emphasis on the proposal adapter for contextual region localization and the seg adapter for pixel-level mask segmentation. To establish a comparative baseline, we design a model configuration that excludes both the proposal and seg adapters, instead relying solely on a switch adapter to provide missing objects and context clues. This baseline model serves as the foundation for evaluating the impact of the individual and combined components of the framework. As demonstrated in Table 5, the introduction of the proposal adapter, the seg adapter, or both, results in progressive and substantial improvements in performance across various evaluation metrics. For instance, accuracy on the $VQA^{v2}$ dataset improves from 62.1% to 63.8%, showcasing the considerable value these components add. This underscores the pivotal roles of the proposal and seg adapters in enhancing the model’s overall capability, further affirming their importance within the FaST framework.
Further, we evaluate the switch adapter’s role in incorporating missing objects and context clues using a variant Baseline*, which omits these features. Table 5 shows that adding missing objects or context clues improves metrics like MME and refCOCOg, with the best performance achieved when both are included. These results confirm the importance of all components in optimizing FaST’s effectiveness.
@@ 3.4 Qualitative Comparisons of FaST
In Fig.6, we present qualitative comparisons that highlight the enhancements introduced by FaST. The top row, above the dotted line, shows results from the VQA task, comparing FaST with LLaVA-v1.5. LLaVA-v1.5 often fails to focus on key areas within the image, leading to incorrect or incomplete responses. In contrast, FaST builds a chain of evidence by identifying key objects and elements (e.g., detecting a woman on the street or a monitor on the roof) and then applying object-level pixel masks to accurately determine the focus areas. This enables FaST to provide more precise and deliberate answers. The bottom row, below the dotted line, shows segmentation results, comparing FaST with LISA-7B. LISA-7B struggles with segmenting smaller objects or those requiring more complex reasoning, often causing confusion. In contrast, FaST excels at isolating relevant objects with greater accuracy and granularity, particularly with smaller or less obvious items. This demonstrates FaST’s superior performance in both VQA and segmentation tasks, showcasing its ability to handle a wide range of visual and reasoning challenges more effectively than its counterparts.
4 Related works
LLM as Visual Agents. With the capabilities that LLMs have demonstrated in language understanding and generation (Ouyang et al., 2022; OpenAI, 2024; Zheng et al., 2023; Touvron et al., 2023a; b; Wang et al., 2024a; Hua et al., 2024; Mei et al., 2024; Jin et al., 2024a), the research community has progressed to explore how LLMs can be enhanced with vision input for multimodal tasks as visual agents (Alayrac et al., 2022; Driess et al., 2023; Li et al., 2023a; Ge et al., 2023; Dai et al., 2024; Liu et al., 2023b; 2024a; Lin et al., 2024). There are two paradigms for LLM-based visual agents: end-to-end based and tool-using visual agents. Following the principle of instruction tuning, end-to-end visual agents are trained with a curated visual instruction tuning dataset to digest features from multi-modality, unlocking the capability to answer visual questions (Huang et al., 2023; Luo et al., 2023; Zhu et al., 2023a; Bai et al., 2023; Zhang et al., 2023b; c; Chen et al., 2024a; Ye et al., 2023; Singh et al., 2019). For other visual tasks ($e.g.$, Segmentation, Detection, etc), end-to-end trained tailored agents can further perform downstream tasks (Pi et al., 2023; Peng et al., 2024; Lai et al., 2024; Chen et al., 2023b; Wang et al., 2023c; Dai et al., 2024; Wang et al., 2024b; 2023b; Jiang et al., 2023; Chen et al., 2023a). Recent research has focused on leveraging improved vision encoders and fostering more detailed visual understanding, yielding promising results (Fan et al., 2024; Xu et al., 2024a; Shi et al., 2024). While these approaches can be implemented with direct instruction tuning data, they represent a ‘System 1’ type of training. This type of training primarily relies on the dataset’s quality and tends to provide direct answers that are prone to hallucinations, a consequence inherent to the nature of System 1 instruction tuning data. For the second paradigm, tool-using models are built on top of a frozen LLM with access to pretrained visual perception tools (Surís et al., 2023; Shen et al., 2023; Lu et al., 2023a). In this scenario, the LLM first selects visual tools and then decides by thoroughly analyzing the fine-grained information extracted by visual tools (Lu et al., 2023a; You et al., 2023; Wu et al., 2024). While external visual tools enhance the interpretability of the reasoning process, their complexity can introduce inaccuracies. Moreover, the abundance of information generated during reasoning may overshadow key details relevant to the query, resulting in incorrect answers.
System 2 in AI. Recently, LLMs have been engineered to produce text that mimics the step-by-step reasoning process characteristic of human cognition, akin to the analytical and deliberate thought processes associated with what is termed as System 2 in the human cognition process (Qiao et al., 2023; Huang & Chang, 2023a; Wang et al., 2023a; Shaikh et al., 2023; Shao et al., 2024). The systematic approach to problem-solving is a hallmark across various domains, including mathematical word problems (Kojima et al., 2022; Wang et al., 2023d; Lightman et al., 2023; Cobbe et al., 2021; Liu et al., 2023c; Zhu et al., 2023b; Lu et al., 2023b), logical reasoning (Yao et al., 2023d; a; Besta et al., 2024; Wen et al., 2023; Lei et al., 2023; Cheng et al., 2024; Jin et al., 2024b), and multi-modal reasoning (Chen et al., 2024d; You et al., 2023; Wu & Xie, 2024). In Explainable AI, this systematic method is emulated by the model as it generates a text-based elucidation of its reasoning and decision-making process through step by step reasoning process ($e.g.$, chain of thought) (Han et al., 2024; Zhao et al., 2024; Jacovi & Goldberg, 2020; Hua & Zhang, 2022). However, it is crucial to recognize that LLMs, while powerful, are not exempt from encountering challenges when facing complex problems. One such challenge is the issue of hallucination (Zhou et al., 2024; Cui et al., 2023; Li et al., 2023b; Zhang et al., 2023a; Chen et al., 2024c; Guan et al., 2024a), which can distort the model’s reasoning process and lead to inaccuracies in the explanations provided. Initially, LLM reasoning is seen as only a linear chain of thoughts, where each step in the reasoning process is clearly articulated. As models evolve, they adopt more complex structures like hierarchical trees (Geng et al., 2023; Yao et al., 2023b) and intricate graphs (Besta et al., 2024), which enable them to handle much more complex problems but also restrict their general applicability because of increased topological complexity (Yao et al., 2023b; a; Besta et al., 2024; Lei et al., 2023; Wen et al., 2023). Moreover, these complex structures can lead to errors propagating through the model’s reasoning, causing a cascade of mistakes (Xu et al., 2024b). To counter this, incorporating feedback from intermediate reasoning steps and employing iterative refinement, which is similar to human reflection, could help mitigate errors (Chu et al., 2023; Tong et al., 2024b; Guan et al., 2024b; Madaan et al., 2023; Yuan et al., 2024; Wu & Xie, 2024). In unsupervised scenarios, such feedback is vital for enhancing the reasoning capabilities of LLMs and reducing errors (Yao et al., 2023c).
Our key contribution is the introduction of the chain of evidence within multimodal reasoning frameworks. This methodology enriches each reasoning step with accurate, image-based cascading information, effectively mirroring human visual and System 2 cognitive processes. Our approach enhances accuracy and significantly improves interpretability and generalization capabilities.